### v2

# Quasi Zigzag Persistence: A Topological Framework for Analyzing Time-Varying Data

Abstract
In this paper, we propose Quasi Zigzag Persistent Homology (QZPH) as a framework for analyzing time-varying data by integrating multiparameter persistence and zigzag persistence. To this end, we introduce a stable topological invariant that captures both static and dynamic features at different scales. We present an algorithm to compute this invariant efficiently. We show that it enhances the machine learning models when applied to tasks such as sleep-stage detection, demonstrating its effectiveness in capturing the evolving patterns in time-varying datasets.
1 Introduction
Time varying data analysis Hamilton (1994); Box & Jenkins (1976) has been a fundamental challenge in the machine learning community for a long time, ranging from traditional univariate and multivariate time-series data to more complex structure such as sequence of graphs or point clouds. While the traditional time-series data has been handled effectively Wu et al. (2020), increasing complexity of modern applications has necessitated novel methodologies. Spatiotemporal Graph Neural Networks Yu et al. (2018); Oreshkin et al. (2021); Kan et al. (2022); Chu et al. (2023) have emerged as powerful tools to analyze graph sequences by decomposing the problem into two components: spatial dependency and temporal evolution. Similarly, specialized architectures have been proposed to process sequences of point clouds Liu et al. (2019); Fan et al. (2021); Huang et al. (2021); Rempe et al. (2020). However, most of these approaches utilize the geometric information, which is mostly local, among other information potentially missing crucial global patterns in the data. This becomes apparent in scenarios where the overall shape and connectivity of the data carry significant meaning, such as in brain connectivity patterns. Topological methods offer a compelling solution by capturing these global structures that persist across different scales. By augmenting the models with topological information, we can develop models that leverage both local geometric patterns and global topological characteristics for improved performance.
Recently, topological Data Analysis (TDA) has emerged as a prominent field that can leverage topological information hidden behind the data into its analysis. Persistent Homology (PH), a cornerstone of TDA, provides a succinct method to extract and represent multiscale topological information. This capability has proved transformative in enhancing machine learning models, particularly where topological properties carry relevant information  Corbet et al. (2019); Chen et al. (2019); Carrière et al. (2020); Kim et al. (2020); Gabrielsson et al. (2020); Hofer et al. (2020); Zhao et al. (2020); Swenson et al. (2020); Carrière & Blumberg (2020); Vipond (2020); Bouritsas et al. (2022); Horn et al. (2022); Cang & Wei (2017); Zhang et al. (2022); Bauer et al. (2023); Liu et al. (2022); Lesnick & Wright (2015). Time-varying data needs extensions of the classical one-parameter persistence homology along two directions. First, the temporal component of the data necessitates another parameter, thus bringing multi parameter persistence homology (MPH) Botnan & Lesnick (2023) into picture. Second, monotone filtrations usually associated with the standard PH computations cannot accommodate deletions that are required for processing time-varying data. This necessitates to use zigzag persistent homology (ZPH) Carlsson & de Silva (2010) which has been shown to capture dynamic topological features in time-series data Myers et al. (2023); Chen et al. (2021); Tinarrage et al. (2025); Hacquard & Lebovici (2024); Coskunuzer et al. (2024); Beltramo et al. (2022). In effect, we need a combination of MPH and ZPH which allows PH (standard filtration) along one parameter and ZPH along the other. This requires changes to the structure of the underlying partially-ordered set (poset) to what we call quasi-zigzag poset resulting into Quasi Zigzag Persistent Homology (QZPH) introducing additional complexity and a new set of challenges.
While MPH captures richer topological information than PH, it presents a different set of challenges due to lack of a complete invariant. This has motivated TDA researchers to look for different invariants that, although not complete, capture sufficient topological information  Vipond (2020)Corbet et al. (2019); Carrière & Blumberg (2020); Russold & Kerber (2024); Scoccola et al. (2024); Loiseaux et al. (2023); Xin et al. (2023); Mukherjee et al. (2024). The juxtaposition of ZPH to MPH needs the zigzag generalizations of these methods.
In this paper, we find that, one of these methods called Gril Xin et al. (2023) that computes a landscape function Bubenik (2015) using generalized ranks of intervals Kim & Mémoli (2021) adapts naturally to the framework of QZPH both conceptually and computationally. This adaptation leads to our key contribution which we call Zz-Gril, a new topological invariant that extends Gril Xin et al. (2023) framework to capture multiscale topological information in time-evolving data. On the theoretical front, we prove the stability of Zz-Gril and show that the generalized rank of a specific type of subposet, in the setting of QZPH, can be computed by an algorithm presented in Dey et al. (2024). We use this to devise an algorithm to compute Zz-Gril, making this method practical for real-world applications. We demonstrate the practical value of Zz-Gril by experimenting on data from critical healthcare monitoring tasks, such as sleep-stage detection via ECG and multivariate time-series data from other applications. We augment the topological information captured by Zz-Gril with existing machine learning models and show that the addition of topological information improves the performance of the models as exhibited by experimental results in Section 5  [^1] Code is available at https://github.com/TDA-Jyamiti/zzgril.
In a very recent paper Flammer & Hüper (2024), the authors consider a similar setup and develop a method to visualize such data using spatiotemporal persistence landscapes. We acknowledge that the setup is similar but there are many differences in the overall results. One main difference is that Flammer & Hüper (2024) uses rectangles as intervals that allow them to use a simple modification of a result in Dey et al. (2024) whereas we use more general subposets for which we prove Theorem 4.5, which allows us to compute generalized ranks with zigzag modules. Because of the same reason, the overall algorithm to compute the landscape also becomes different. Additionally, we give an efficient algorithm to build a quasi zigzag bi-filtration from an input time-varying data. Finally, we develop a complete ML pipeline that we experiment with real world time-varying data which is not considered in Flammer & Hüper (2024).
2 Overview
We begin by recalling the definition of a simplicial complex. Given a finite vertex set $V$, a simplicial complex $K=K(V)$ is a collection of subsets of $V$ such that if a subset $\sigma\subseteq V$ is in $K$, then all subsets $\tau\subset\sigma$ are also in $K$. Each element in $K$ with cardinality $k+1$ is called a $k$-simplex or simply a simplex. A (non-zigzag) filtration $\mathcal{F}$ is a nested sequence of simplicial complexes indexed by integers/natural numbers: $\mathcal{F}:K_{0}\hookrightarrow K_{1}\hookrightarrow\cdots\hookrightarrow K_{n}$ where the inclusion arrows are all in forward direction. If these arrows are not all in forward direction, we get a filtration called zigzag filtration
$\mathcal{Z}:K_{0}\hookrightarrow K_{1}\hookleftarrow\cdots\hookrightarrow K_{n}.$
Extending the indexing to two parameters such as a grid of integers $\mathbb{Z}^{2}$, we get a bi-filtration (possibly zigzag). We will be interested in a bi-filtration where the filtration in one direction, say $y$-direction, is non-zigzag and the filtration in the other, $x$-direction, is zigzag. For a moving point cloud data (PCD) such filtrations arise naturally. For example, consider the case in Figure 1. The moving PCD with 3 points are shown at the bottom row. Along the $y$-direction, we increase the threshold $\delta>0$ so that any two points are joined by an edge if the distance between them is no more than $\delta$. Then, along $y$-direction, for every PCD, we get a non-zigzag filtration whereas along $x$-direction, for every fixed $\delta$, we get a zigzag filtration. We call this mixing of zigzag and non-zigzag a quasi zigzag bi-filtration.
Figure 1: (left) A PCD (bottom row) and a resulting quasi zigzag bi-filtration; (middle) corresponding quasi zigzag persistence module and three intervals (light blue, yellow, green); (right) a rectangle (light yellow) is expanded for obtaining landscape width.
From such a quasi zigzag bi-filtration, we obtain a so-called persistence module by considering the simplicial homology group $H_{p}(K_{i})$ in some dimension $p$ for every simplicial complex $K_{i}$. The inclusion maps in zigzag or non-zigzag filtrations translate to homomorphisms among these homology groups which become linear maps when homology groups turn into vector spaces under a field coefficient, say $k$. These vector spaces with dimension, say $d$, are often depicted with their isomorphic counterparts of $d$ copies of $k$, written as $k^{d}$. We refer to such a persistence module as a quasi zigzag persistence module; In Figure 1 (right), we show the quasi zigzag persistence module for zeroth homology group $H_{0}(\cdot)$ that represents the number of components.
The indexing set of a non-zigzag, zigzag, or quasi zigzag filtration can be thought of as a poset $P$ where for every pair of comparables $\mathbf{p}\leq\mathbf{q}$ in $P$, one has the inclusion of complexes $K_{\mathbf{p}}\subseteq K_{\mathbf{q}}$. A non-zigzag poset can be indexed by integers in $\mathbb{Z}$. A zigzag poset that arises in our case is a subposet of the poset $\mathbb{ZZ}$ defined below.
Definition 2.1 ((Quasi) Zigzag Poset).
Let $\mathbb{ZZ}$ be the subposet of $\mathbb{Z}^{\mathsf{op}}\times\mathbb{Z}$ defined as
$\mathbb{ZZ}\coloneqq\{(i,j):i\in\mathbb{Z}^{\mathsf{op}},j\in\{i,i-1\}\}.$
A poset $P$ is called zigzag if $P$ is a subposet of $\mathbb{ZZ}$ and is quasi zigzag if it is a subposet of the product $\mathbb{ZZ}\times\mathbb{Z}$ with the product order.
Now, we formally define persistence modules over different kind of posets.
Definition 2.2 (Persistence module).
A persistence module $M$ over a poset $P$ is a functor $M:P\rightarrow{\bf vec}$ where $P$ is considered as a category and ${\bf vec}$ is the category of finite dimensional vector spaces. Stating more explicitly, $M$ is a collection of vector spaces $\{M_{\mathbf{p}}\}_{\mathbf{p}\in P}$ and linear maps $\{M_{\mathbf{p}\rightarrow\mathbf{q}}\}$ for every comparable $\mathbf{p}\leq\mathbf{q}$ in $P$. It is called a zigzag module if $P$ is zigzag and a quasi zigzag module if $P$ is quasi zigzag.
Next, we introduce the concepts of connectedness and intervals in posets which will be used for further exposition. We say a poset $P$ is connected if there is a sequence of points $p=r_{1},\ldots,r_{t}=q$ in $P$ so that $r_{i}\leq r_{i+1}$ or $r_{i+1}\leq r_{i}$ for every $i\in\{1,\ldots,t-1\}$. We say a subposet $I\subseteq P$ is convex if for every pair $p,q\in I$ and every $p\leq r\leq q$, we have $r$ in $I$. An interval $I$ in a poset $P$ is a subposet which is both convex and connected. An interval $I$ with $m\geq 0$ minima and $n\geq 0$ maxima is called an $(m,n)$-interval. In the special case of $\mathbb{Z}^{2}$, every interval has a special form where the sets of minima and maxima along with other points constitute staircases that can be visualized as part of the boundary of the interval; see Figure 2.
The main construct upon which we build Zz-Gril involves the so-called generalized ranks  Kim & Mémoli (2021) which can be computed efficiently by a zigzag persistence algorithm Dey et al. (2024). Let $P$ be a quasi zigzag poset over which a quasi zigzag persistence module $M$ is defined and $I\subseteq P$ be a subposet. When $M$ is induced by a bi-filtration in QZPH, the generalized rank of the module restricted to $I$ measures the multiplicity of the homological features (# of independent homological classes) that have support over the entire poset $I$. For example, considering $I=P$ (shaded light green), we see that, in Figure 1(middle), one component survives the entire poset. Then, its generalized rank for zero-th homology $H_{0}$ is 1. Similarly, the subposet in the bottom row (light blue) has a generalized rank of 3 for $H_{0}$ and the subposet in bottom right square (light yellow) has a generalized rank of 2 for $H_{0}$. We formally define generalized rank below.
Definition 2.3 (Generalized rank).
Let $M:P\to\mathbf{vec}$ be a persistence module from a finite and connected poset $P$. The restriction of $M$ to a subposet $I$ of $P$, $M|_{I}$, is the collection of vector spaces $M_{\mathbf{p}}$ for $\mathbf{p}\in I$ along with linear maps $M_{\mathbf{p}\to\mathbf{q}}$ for all comparable $\mathbf{p}\leq\mathbf{q}$ in $I$. Then, the generalized rank of $M$ over $I$ is defined as the rank of the canonical linear map from $\mathsf{lim}M|_{I}$ to $\mathsf{colim}M|_{I}$
$\mathsf{rk}^{M}(I)\coloneqq\text{rank}\left(\mathsf{lim}M|_{I}\to\mathsf{colim}M|_{I}\right).$
We refer the reader to MacLane (1971) for the definitions of limit, colimit, and the construction of the canonical limit-to-colimit map.
One of the very useful properties of generalized rank is that it is monotone w.r.t. the inclusion of posets. We use this property to define Zz-Gril.
Monotone Property of Generalized Rank. It follows from the definition that $\mathsf{rk}^{M}(I)\leq\mathsf{rk}^{M}(J)$ for all $J\subseteq I$, where $I\text{ and }J$ are subposets in $P$.
To exploit the topological features captured by generalized ranks, we use a cover of the quasi zigzag poset with a specific set of subposets called worms. When endowed with $\mathbb{Z}^{2}$ ordering, these subposets become $(3,3)$-intervals around a set of chosen centers with some chosen widths. Definition 3.1 gives precise definitions of these terms. We expand each of these worms, that is, increase the widths of these $(3,3)$-intervals while keeping the centers fixed. The monotonicity of the generalized rank ensures that it can only decrease or remain the same by this expansion. In Figure 1(right) a $(1,1)$-interval (rectangle) is expanded maximally to decrease the generalized rank from $3$ to $1$. The width for which the rank drops below a chosen threshold is taken as the Zz-Gril value at the center of the interval. Defintion 3.2 makes this concept precise. The set of Zz-Gril values at the chosen centers makes a vector that we use in the learning pipeline.
3 Zz-Gril
In this section, we introduce concepts and definitions required for Zz-Gril. Then, we define Zz-Gril and discuss its theoretical properties. Keep in mind that, even if we do not explicitly state it always, these concepts apply to QZPH setting where vector spaces and linear maps in question are given by simplicial homology groups and homomorphisms between them induced by the input bi-filtrations.
Consider the poset $\mathbb{ZZ}\times\mathbb{Z}$ with the product order, i.e., $(\mathbf{z_{1}},z_{2})\leq(\mathbf{w_{1}},w_{2})$ if $\mathbf{z_{1}}\leq\mathbf{w_{1}}$ in $\mathbb{ZZ}$ and $z_{2}\leq w_{2}$ in $\mathbb{Z}$. Note that $\mathbb{ZZ}\times\mathbb{Z}$ is equivalent to $\mathbb{Z}^{2}$ as sets. Thus, every subposet $P\subseteq\mathbb{ZZ}\times\mathbb{Z}$ can be thought as a subposet $P^{\mathbb{Z}^{2}}\subseteq\mathbb{Z}^{2}$ that is endowed with $\mathbb{Z}^{2}$ ordering. Interestingly, an interval in $\mathbb{Z}^{2}$ may not remain an interval in $\mathbb{ZZ}\times\mathbb{Z}$ and vice-versa.
We define a quasi zigzag bi-filtration over a quasi zigzag poset $P$ as a collection of simplicial complexes $\{K_{\mathbf{p}}\}_{\mathbf{p}\in P}$, where $K_{\mathbf{p}}\subseteq K_{\mathbf{q}}$ for all comparable $\mathbf{p}\leq\mathbf{q}$. Akin to 1-parameter and zigzag persistence modules, inclusion-induced linear maps between the homology vector spaces of simplicial complexes in a quasi zigzag bi-filtration define a quasi zigzag persistence module.
We now define a special type of subposet in $\mathbb{ZZ}\times\mathbb{Z}$ which in $\mathbb{Z}^{2}$ is an $\ell$-worm introduced by Xin et al. (2023) for $\ell=2$.
Definition 3.1 (Worm).
Let $\mathbf{p}\in\mathbb{ZZ}\times\mathbb{Z}$ and $\delta\in\mathbb{Z}$ be given. Let $\boxed{\mathbf{p}}_{\delta}$ denote the $\delta$-square centered at $\mathbf{p}$, i.e., $\boxed{\mathbf{p}}_{\delta}\coloneqq\{\mathbf{z}\in\mathbb{ZZ}\times\mathbb{Z}:||\mathbf{p}-\mathbf{z}||_{\infty}\leq\delta\}$. Then, a worm centered at $\mathbf{p}$ with width $\delta$ is defined as the union of $\boxed{\mathbf{p}}_{\delta}$ with the two $\delta$-squares $\boxed{\mathbf{q}}_{\delta}$ centered at points $\mathbf{q}=\mathbf{p}\pm(\delta,-\delta)$ on the off-diagonal line segment. We denote the worm as $\boxed{\mathbf{p}}_{\delta}^{2}$. The superscript denotes the number of $\delta$-squares in the union apart from $\boxed{\mathbf{p}}_{\delta}$.
We note that computing $||\cdot||_{\infty}$ in $\mathbb{ZZ}\times\mathbb{Z}$ is equivalent to computing it in $\mathbb{Z}^{2}$ using the notion of set equivalence. We can see that a worm turns out to be a $(3,3)$-interval in $\mathbb{Z}^{2}$. Refer to Figure 2 for an illustration of a worm.
Definition 3.2 (Zz-Gril).
Let $M$ be a quasi zigzag persistence module. Then, the ZigZag Generalized Rank Invariant Landscape is a function $\lambda^{M}:\mathbb{ZZ}\times\mathbb{Z}\times\mathbb{N}\to\mathbb{N}$ defined as
$\lambda^{M}(\mathbf{p},k)\coloneqq\sup\left\{\delta\geq 0:\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq k\right\},$
where $\mathbf{p}\in\mathbb{ZZ}\times\mathbb{Z}$.
The basic idea of Zz-Gril is similar to Gril. However, we note that the underlying poset structure is very different. We cover $\mathbb{ZZ}\times\mathbb{Z}$ with subposets of specific shapes (worms), and compute generalized rank over these worms to define the landscape function (Zz-Gril). Next, we show that Zz-Gril as an invariant remains stable.
Figure 2: The worm with blue boundary represents a worm centered at $\mathbf{p}$ with width $\delta=1$. The worm with red boundary represents an expanded worm, centered at $\mathbf{p}$ with width $\delta=2$.
3.1 Stability of Zz-Gril
We prove the stability of Zz-Gril by showing that its perturbation is bounded by the interleaving distance between two quasi zigzag persistence modules. The definition of interleaving distance (see Definition A.3 in Appendix A) between two zigzag persistence modules can be extended to quasi zigzag persistence modules.
There is an alternate notion of proximity on the space of persistence modules which uses Generalized Ranks computed on all intervals. This is known as erosion distance Kim & Mémoli (2021). We define a distance similar to erosion distance on the space of quasi zigzag persistence modules based on their generalized ranks computed over worms. For this, we need the notion of $\epsilon$-thickening.
Let $(\mathbb{ZZ}\times\mathbb{Z})$ denote the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals.
Definition 3.3.
Let $\mathcal{L}$ denote the collection of all worms in $\mathbb{ZZ}\times\mathbb{Z}$. Let $M$ and $N$ be quasi zigzag persistence modules. The erosion distance is defined as:
$d_{\mathcal{E}}^{\mathcal{L}}(M,N)\coloneqq\inf\left\{\epsilon\geq 0:\forall\boxed{\mathbf{p}}^{2}_{\delta}\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\text{ and }\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\right\}.$
Proposition 3.4.
Given two quasi zigzag persistence modules $M$ and $N$, $d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{I}}(M,N)$ where $d_{\mathcal{I}}$ denotes the interleaving distance between $M$ and $N$.
Proposition 3.4 leads us to Theorem 3.5.
Theorem 3.5.
Let $M$ and $N$ be two quasi zigzag persistence modules. Given appropriate $\mathbf{p},k$, let the Zz-Gril of $M$ and $N$ over $\boxed{\mathbf{p}}^{2}_{\delta}$ be $\lambda^{M}$ and $\lambda^{N}$ respectively. Then,
$|\lambda^{M}(\mathbf{p},k)-\lambda^{N}(\mathbf{p},k)|=d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{I}}(M,N).$
Corollary 3.6 (Stability).
Let $M$ and $N$ be two quasi zigzag persistence modules. Then,
$||\lambda^{M}-\lambda^{N}||_{\infty}\leq d_{\mathcal{I}}(M,N).$
We refer the reader to Appendix A for the proofs of Proposition 3.4 and Theorem 3.5.
4 Algorithm
We discuss the details of computing Zz-Gril in this section. We begin by proving that generalized rank over a subposet in $\mathbb{ZZ}\times\mathbb{Z}$ can also be computed by computing the number of full bars in the zigzag filtration along the “boundary” of the subposet. To be precise, we introduce the following concepts now. Let $P$ be any finite connected poset. We say a point $q\in P$ covers a point $p\in P$, denoted $p\prec q$ or $q\succ p$, if $p\leq q$ and there is no $r\in P$, $r\not=p,q$, so that $p\leq r\leq q$. A covering path $p\sim q$ between $p$ and $q$ is either a set of points $\{p=p_{0}\prec p_{1}\prec\cdots\prec p_{\ell}=q\}$ or $\{p=p_{0}\succ p_{1}\succ\cdots\succ p_{\ell}=q\}$. We define some special points in $P$. The set of minima in $P$ are the points $P_{min}=\{p\in P\ \,|\,\not\exists r\not=p\in P\mbox{ where }r\leq p\}$. Similarly, define the set of maxima $P_{max}$. For any two points $p,q\in P$, let $p\vee q$ and $p\wedge q$ denote their least upper bound (lub) and largest lower bound (llb) in $P$ if they exist.
Now consider a subposet $I\subseteq\mathbb{ZZ}\times\mathbb{Z}$ whose counterpart $I^{\mathbb{Z}^{2}}$ with $\mathbb{Z}^{2}$ ordering becomes an interval. Since $I^{\mathbb{Z}^{2}}\subset\mathbb{Z}^{2}$, its points can be sorted with increasing $x$-coordinates in $\mathbb{Z}^{2}$. Let $I^{\mathbb{Z}^{2}}_{min}=\{p_{0},\ldots,p_{s}\}$ and $I^{\mathbb{Z}^{2}}_{max}=\{q_{0},\ldots,q_{t}\}$ be the set of minima and maxima respectively sorted according to their $x$-coordinates. The lower fence and upper fence of $I^{\mathbb{Z}^{2}}$ are defined as the paths
$L_{I}^{\mathbb{Z}^{2}}:=p_{0}\sim(p_{0}\vee p_{1})\sim p_{1}\sim(p_{1}\vee p_{2})\sim p_{2}\sim\cdots\sim p_{s}$
$U_{I}^{\mathbb{Z}^{2}}:=q_{0}\sim(q_{0}\wedge q_{1})\sim q_{1}\sim(q_{1}\wedge q_{2})\sim q_{2}\sim\cdots\sim q_{t}.$
Consider the boundary $B_{I}^{\mathbb{Z}^{2}}$ comprising $L_{I}^{\mathbb{Z}^{2}}$, $U_{I}^{\mathbb{Z}^{2}}$, the paths $p_{0}\sim q_{0}$ and $p_{s}\sim q_{t}$ going through the top left and bottom right corner points of $I^{\mathbb{Z}^{2}}$ respectively. Consider the subposet $I\subseteq\mathbb{ZZ}\times\mathbb{Z}$ and observe that $I_{min},I_{max}\subseteq B_{I}^{\mathbb{Z}^{2}}$ because all other points have one point below and another above in the $y$-direction. Let $\partial_{L}I$, $\partial_{U}I$, and $\partial I$ be minimal paths in $BI^{\mathbb{Z}^{2}}$ connecting all minima in $I_{min}$, all maxima in $I_{max}$, and all minima, maxima together respectively. Drawing an analogy to Dey et al. (2024) we call $\partial I$ a boundary cap which is drawn orange in Figure 3.
Next, we appeal to certain results in category theory to claim a result analogous to Dey et al. (2024) that helps computing the generalized ranks with zigzag persistence modules. Treating $I$ as a category with points in $I$ as objects and relations as morphisms, we get $\partial_{L}I$ and $\partial_{U}I$ as subcategories. Then, we get functors $F_{L}:\partial_{L}I\rightarrow I$ induced by inclusion and $F_{U}:I\rightarrow\partial_{U}I$ induced by projection which have some special properties.
Definition 4.1.
A connected subposet $P^{\prime}\subseteq P$ is called initial if for every $p\in P\setminus P^{\prime}$, the downset $\downarrow{p}=\{q\in P\,|\,q\neq p~{}\&~{}q\leq p\}$ intersects $P^{\prime}$ in a connected poset. Similalrly, $P^{\prime}$ is called terminal if for every $p\in P\setminus P^{\prime}$, the upset $\uparrow{p}=\{q\in P\,|\,q\neq p~{}\&~{}q\geq p\}$ intersects $P^{\prime}$ in a connected poset.
We call a functor $F:P^{\prime}\rightarrow P$ for $P^{\prime}\subseteq P$ initial if $P^{\prime}$ is initial. Similarly, we call $F:P\rightarrow P^{\prime}$ terminal if $P^{\prime}$ is terminal. The following result connecting initial (terminal) functor to the limit (colimit) is well known, [Chapter 8] Riehl (2014).
Proposition 4.2.
Let $M:P\rightarrow{\bf vec}$ be a persistence module and $F_{L}:P^{\prime}\rightarrow P$ and $F_{U}:P\rightarrow P^{\prime\prime}$ be initial and terminal functors. Then, there is an isomorphism $\phi:\mathsf{lim}{M}\rightarrow\mathsf{lim}{M|_{P^{\prime}}}$ from the limit of $M$ to the limit of the restricted module $M|_{P^{\prime}}$. Similarly, we have an isomorphism for colimits, $\psi:\mathsf{colim}{M|_{P^{\prime\prime}}}\rightarrow\mathsf{colim}{M}$.
We observe that the functors $F_{L}:\partial_{L}I\rightarrow I$ and $F_{U}:I\rightarrow\partial_{U}I$ mentioned earlier are initial and terminal respectively allowing us to use Proposition 4.2.
Proposition 4.3.
$F_{L}:\partial_{L}I\rightarrow I$ is an initial functor and $F_{U}:I\rightarrow\partial_{U}I$ is a terminal functor.
Proof.
It is sufficient to prove that $\partial_{L}I$ is initial for $F_{L}$ and $\partial_{U}I$ is terminal for $F_{U}$. We only show that $\partial_{L}I$ is initial and the proof for $\partial_{U}I$ being terminal is similar.
Let $p\in I\setminus\partial_{L}I$ be any point. Let $p^{-}$ and $p^{+}$ be two points (if exist) where $p^{-}\prec p$ and $p^{+}\succ p$ in $\mathbb{Z}^{2}$ with $p^{-}_{x}<p_{x}<p^{+}_{x}$. By definition of the quasi zigzag poset, either $p^{-}>p$ and $p^{+}>p$, or $p^{-}<p$ and $p^{+}<p$ in the poset $\mathbb{ZZ}\times\mathbb{Z}$.
In the first case, the downset $\downarrow p$ consists of all points $p^{\prime}\neq p$ that are vertically below $p$, that is, $p^{\prime}_{x}=p_{x}$ and $p^{\prime}_{y}<p_{y}$. This means $\downarrow p$ intersects $\partial_{L}I$ in a connected poset $p^{\prime}_{0}<\cdots<p^{\prime}_{k}$ where each $p^{\prime}_{i}$ has the same $x$-coordinate as $p$.
In the second case, the downset $\downarrow p$ consists of three sets of points $Y^{p^{-}}$, $Y^{p}$, and $Y^{p^{+}}$ that are vertically below $p^{-}$, $p$, and $p^{+}$ respectively. Each of $Y^{p^{-}}$, $Y^{p}$ and $Y^{p^{+}}$ intersects $\partial_{L}I$ in a connected poset. They can be disconnected as a whole only if there is a point $q\in\partial_{L}I$ so that $p^{-}_{x}<q_{x}<p_{x}$ or $p_{x}<q_{x}<p^{+}_{x}$. But, neither is possible because $p^{-}$ and $p^{+}$ are points in $I$ covered by $p$. ∎
Following proposition is immediate.
Proposition 4.4.
$\partial(\partial_{L}I)=\partial_{L}I$ and $\partial(\partial_{U}I)=\partial_{U}I$.
Now, we prove the following result which extends Theorem 3.1 in Dey et al. (2024) from $2$-parameter persistence modules to the quasi zigzag persistence modules.
Theorem 4.5.
Let $M:P\rightarrow{\bf vec}$ be a quasi zigzag persistence module and $I\subseteq P$ be a finite subposet so that $I^{\mathbb{Z}^{2}}$ is an interval. Then, $\mathsf{rk}^{M}(I)=\mathsf{rk}^{M}(\partial I)$.
Proof.
Let $r$ denote the map $r:\mathsf{lim}{I}\rightarrow\mathsf{colim}{I}$, and $\phi:\mathsf{lim}{I}\rightarrow\mathsf{lim}{\partial_{L}I}$ and $\psi:\mathsf{colim}{\partial_{U}I}\rightarrow\mathsf{colim}{I}$ be the isomorphisms guaranteed by Proposition 4.2. Observe that $\mathsf{rk}^{M}(I)=\mathrm{rank}(\psi^{-1}\circ r\circ\phi^{-1})$. Now let $r^{\prime}$ denote the map $r^{\prime}:\mathsf{lim}{\partial I}\rightarrow\mathsf{colim}{\partial I}$. Consider the isomoprhisms $\phi^{\prime}:\mathsf{lim}{\partial I}\rightarrow\mathsf{lim}{\partial(\partial_{L}I)}$ and $\psi^{\prime}:\mathsf{colim}{\partial(\partial_{U}I)}\rightarrow\mathsf{colim}{\partial I}$ again guaranteed by Proposition 4.2. Then, $\mathsf{rk}^{M}(\partial I)=\mathrm{rank}(\psi^{\prime-1}\circ r^{\prime}\circ\phi^{\prime-1})$. But, $\psi^{-1}\circ r\circ\phi^{-1}=\psi^{\prime-1}\circ r^{\prime}\circ\phi^{\prime-1}$ because of Proposition 4.4. ∎
With this background, the pipeline for computing Zz-Gril in the QZPH framework can be described as follows. Suppose that we are given sequential data (sequence of point clouds, sequence of graphs, multivariate time series) with vertex-level correspondences between consecutive time steps. First, we build a quasi zigzag bi-filtration out of this data where time increases in $x$-direction and the threshold for constructing complexes increases in $y$-direction. The algorithm for building this bi-filtration out of raw data is described in section 4.1 where we make certain choices to make it efficient. Each simplicial complex in the bi-filtration is indexed by a finite grid $G$ in $\mathbb{Z}^{2}$ because $\mathbb{ZZ}\times\mathbb{Z}$ is equivalent to $\mathbb{Z}^{2}$ as sets. We sample a set of center points $S$ from $G$ and compute Zz-Gril (Definition 3.2) at each of these center points. Taking advantage of Theorem 4.5, we compute Zz-Gril by computing the zigzag filtration along the boundary cap (a path) of a worm centering each point in $S$ and then computing the number of full bars in the corresponding zigzag persistence module obtained by applying the homology functor.
Figure 3: A worm $I$ (shaded) on a quasi zigzag bi-filtration where the direction of the inclusion on horizontal arrows are shown at the bottom and on the vertical arrows on left. The worm centered at (3,3) has width 1. The part of the boundary colored green is $\partial_{L}I$ connecting the minima shown as green points and the part in purple is $\partial_{U}I$ connecting the maxima shown as purple points. The boundary cap is shown in orange which runs parallel to the boundary for a portion of it. Refer to Figure 5 for the zigzag filtration along the boundary cap of the worm. A sequence of graphs with $T=3$ time steps is shown with the corresponding graph filtrations: $\mathcal{F}_{t_{1}},\mathcal{F}_{t_{2}},\mathcal{F}_{t_{3}}$ each with $L=5$ levels. The filtration of the unions are encircled by ovals. Zigzag filtration at the topmost level $\mathcal{Z}_{L}$ is shown in a rectangular box.
A sequence of point clouds (Figure 1) or graphs (Figure 3) or multivariate time-series data (refer section 5 for converting multivariate time-series to a sequence of point clouds or graphs) gives us a collection of simplicial complexes. Every simplex in each simplicial complex is assigned a weight which is derived from the input. We build a quasi zigzag bi-filtration from this collection of simplicial complexes.
4.1 Building Quasi Zigzag Bi-filtration
We explain the algorithm by considering an example of sequential graph data, as shown in Figure 4. We assume that each graph has edge-weights. In the top row of Figure 4, notice that the first two graphs can not be linked by an inclusion in any direction, i.e., neither is a subgraph of the other. This is because, there is an inclusion of an edge $(v_{2},v_{3})$ as well as a deletion $(v_{1},v_{4})$. We circumvent this problem by clubbing all inclusions together followed by all deletions. This is equivalent to considering the union of the two graphs as the intermediary step and then deleting the edges which are not present in the previous graph. Refer to the bottom row of Figure 4, where the intermediary union graphs are shown along with the original graphs.
A sequence of $T$ number of graphs is converted into a zigzag filtration $\mathcal{Z}_{L}$ of length $12T-$ 1 by the above procedure because every consecutive pair of graphs introduces a union in between. For each graph in the zigzag filtration thus obtained, we construct its graph filtration based on edge-weights; see e.g. Dey & Wang (2022); Edelsbrunner & Harer (2010). We ensure that the length of each graph filtration remains the same by considering the sublevel sets of exactly $L$ levels. The zigzag filtration at the highest level, $\mathcal{Z}_{L}$, can be pulled back to each lower level $l$. This ensures that we get a zigzag filtration $\mathcal{Z}_{l}$ at each level $1\leq l<L$ of the filtration. This gives us a quasi zigzag bi-filtration. Refer to Figure 3 for an illustation of a quasi zigzag bi-filtration.
While implementing this procedure, we need not compute the union graphs explicitly saving both storage and time because all necessary information is already present in the corresponding component graphs. We use the following procedure to efficiently build the quasi zigzag bi-filtration.
Construct the graph filtration $\mathcal{F}_{t_{i}}$ of each of the $T$ graphs in the sequential graph data, where $1\leq t_{i}\leq T$. Let $\sigma$ be a simplex that is inserted at the level $l$ in the filtration $\mathcal{F}_{t_{i}}$. We have three possible scenarios for $\sigma$ in $\mathcal{F}_{t_{i+1}}$:
  1. 1.
    $\sigma$ is inserted in $\mathcal{F}_{t_{i+1}}$ at $m$ for some $m<l$: In this case, $\sigma$ needs to be added on all the horizontal inclusion arrows $K_{t_{i},w}\xhookrightarrow{}K_{t_{i},w}\cup K_{t_{i+1},w}$ for $m\leqslant w<l$.
  2. 2.
    $\sigma$ is inserted in $\mathcal{F}_{t_{i+1}}$ at $m$ for some $m\geqslant l$: In this case, $\sigma$ needs to be added on all the horizontal inclusion arrows $K_{t_{i+1},w}\xhookrightarrow{}K_{t_{i},w}\cup K_{t_{i+1},w}$ for $l\leqslant w<m$.
  3. 3.
    $\sigma$ is not present in $\mathcal{F}_{t_{i+1}}$. In this case, we treat it as getting inserted at the maximum level $L$ in $\mathcal{F}_{t_{i+1}}$ and hence, will be added by the inclusion $K_{t_{i+1},L}\xhookrightarrow{}K_{t_{i},L}\cup K_{t_{i+1},L}$.
We repeat this procedure for the pair $\mathcal{F}_{t_{i}}$ and $\mathcal{F}_{t_{i-1}}$.
Figure 4: The top row shows a typical input to the Zz-Gril framework as a sequence of graphs. This sequence has 3 graphs. Observe that consecutive graphs cannot be linked by an inclusion in either direction because neither is a subgraph of the other. In order to circumvent this problem, we consider the unions of consecutive graphs as an intermediary step, as shown in the bottom row.
Since we are working with an integer grid, we can simulate the quasi zigzag bi-filtration such that we have the standard filtration $\mathcal{F}_{t_{i}}$ at even $x$-coordinates and at odd $x$-coordinates, we have the standard filtration of the unions, without explicitly storing the unions.
Correctness. To prove the correctness of the procedure, we need to show the following: For $1\leq i,i^{\prime}\leq T$ and $1\leq j,j^{\prime}\leq L$, let $K_{i,j}$ denote the simplicial complex in the quasi zigzag filtration built according to the procedure mentioned above and $K^{\prime}_{i,j}$ is the simplicial complex in the quasi zigzag bi-filtration built by considering unions explicitly. We show that the new simplices inserted by the inclusion $K_{i,j}\xhookrightarrow{}K_{i^{\prime},j^{\prime}}$ are the same as the ones inserted by $K^{\prime}_{i,j}\xhookrightarrow{}K^{\prime}_{i^{\prime},j^{\prime}}$.
Proof.
The statement is immediately true for the inclusions $i^{\prime}=i,j^{\prime}=j+1$. This is because the new simplices in the upward ($y$-direction) inclusions are the ones born at the level $j+1$ in the input family of filtrations. Now, we prove the statement for the case where $i^{\prime}=i\pm 1,j^{\prime}=j$, i.e., the horizontal ($x$-direction) inclusions/deletions. Note that we have inclusions from $K_{i,j}\xhookrightarrow{}K_{i^{\prime},j^{\prime}}$ only when $i$ is an even integer and $i^{\prime}=i\pm 1$. We prove it for the case $i^{\prime}=i+1,j^{\prime}=j$. The proof for the other case is similar. Let $\sigma$ be a simplex added on the inclusion $K_{i,j}\xhookrightarrow{}K_{i+1,j}$. According to the procedure described above, it means that $\sigma$ is born at some level $j_{1}\leq j$ in the filtration $\mathcal{F}_{i+2}$ and is not born before level $j$ in the filtration $\mathcal{F}_{i}$. Hence, $\sigma$ will be newly added on the arrow $K^{\prime}_{i,j}\xhookrightarrow{}K^{\prime}_{i+1,j}$ because $\sigma\in K^{\prime}_{i+2,j}$ and $\sigma\notin K^{\prime}_{i,j}$. This is because $K_{i,j}=K^{\prime}_{i,j}$ for all even integers $i$, as they are the simplicial complexes part of the input filtrations $\mathcal{F}_{i}$. By the same arguments, all the newly added simplices on inclusions $K^{\prime}_{i,j}\xhookrightarrow{}K^{\prime}_{i+1,j}$ will be newly added simplices on $K_{i,j}\xhookrightarrow{}K_{i+1,j}$. ∎
Figure 5: This is the zigzag filtration $\mathcal{Z}_{bdry}$ along the boundary cap of the worm shown in Figure 3.
Figure 6: Pipeline used for the experiments.
4.2 Computing Zz-Gril
Given this quasi zigzag bi-filtration, we obtain a quasi zigzag persistence module $M$ by considering the homology $\{H_{p}(K_{t_{i},\alpha_{j}})\}_{i,j}$ of each simplicial complex. Given a sampled center point $\mathbf{p}$ and the value of rank $k$, we need to compute $\lambda^{M}(\mathbf{p},k)$. Thus, we need to compute the maximal width of the worm such that the value of generalized rank of $M$ over the worm is at least $k$. We do a binary search over all the possible widths of the worms to arrive at such maximal width for each tuple $(\mathbf{p},k)$. To compute the generalized rank of $M$ over a worm, we need the zigzag filtration along the boundary cap of the worm, $\mathcal{Z}_{bdry}$. We utilize the stored information about which simplices to add/delete along the horizontal and vertical arrows in order to build $\mathcal{Z}_{bdry}$. Then, we compute the number of full bars in the zigzag persistence module corresponding to $\mathcal{Z}_{bdry}$ using an efficient zigzag algorithm proposed in Dey & Hou (2022) and use it in the binary search routine. The pseudocode of this idea is shown in Algorithm 1.
Algorithm 1 ComputeZz-Gril
  Input: $\mathcal{ZZ}:\text{Quasi zigzag bi-filtration},k\geq 1,\mathbf{p}$
  Output: $\lambda(\mathbf{p},k)$: Zz-Gril value at point $\mathbf{p}$ for fixed $k$
  Initialize: $\delta_{min}\leftarrow 1,\delta_{max}\leftarrow\text{len}(\mathcal{F})$, $\lambda\leftarrow 1$
  while $\delta_{min}\leq\delta_{max}$ do
     $\delta\leftarrow(\delta_{min}+\delta_{max})/2$
     $I\leftarrow\boxed{\mathbf{p}}^{2}_{\delta}$;   $r\leftarrow$ ComputeRank($\mathcal{ZZ},I$)
     if $r\geq k$ then
        $\lambda\leftarrow\delta$;   $\delta_{min}\leftarrow\delta+1$
     else
        $\delta_{max}\leftarrow\delta-1$
     end if
  end while
  return $\lambda$
5 Experiments
In this section, we report the results of testing Zz-Gril on various datasets. We begin by giving a detailed description about the experimental setup. Then, we give a brief description of the datasets, followed by the experimental results. We use the benchmark UEA multivariate time-series Bagnall et al. (2018) datasets to test Zz-Gril on multivariate time series data to show that Zz-Gril can be applied to datasets from various domains. Further, we test Zz-Gril on a targeted application of sleep-stage classification by performing experiments on ISRUC-S3 Khalighi et al. (2016) dataset. In all these experiments, we augment the topological information captured by Zz-Gril to one of the specifically tailored machine learning methods on the respective datasets and compare.
5.1 Experimental Setup
Each data instance is a multivariate time-series which we convert into a quasi zigzag bi-filtration. The Zz-Gril framework takes in a quasi zigzag bi-filtration as input and provides a topological signature for the sequence as output. We augment the machine learning model with this topological information and train the model for classification. The framework is shown in Figure 6.
Dataset/Methods ED-1NN DTW-1NN-I DTW-1NN-D MLSTM-FCN ShapeNet WEASEL+MUSE OS-CNN MOS-CNN Zz-Gril
FingerMovements 0.550 0.520 0.530 0.580 0.589 0.490 0.568 0.568 0.590
Heartbeat 0.620 0.659 0.717 0.663 0.756 0.727 0.489 0.604 0.721
MotorImagery 0.510 0.390 0.500 0.510 0.610 0.500 0.535 0.515 0.580
NATOPS 0.860 0.850 0.883 0.889 0.883 0.870 0.968 0.951 0.850
SelfRegulationSCP2 0.483 0.533 0.539 0.472 0.578 0.460 0.532 0.510 0.522
Table 1: Acccuracy comparison of Zz-Gril with some existing methods on UEA Multivariate Time Series Classification Datasets. Bold entries denote the best model and gray the second best.
Dataset/Methods TapNet TapNet+Zz-Gril TodyNet TodyNet+Zz-Gril
FingerMovements 0.530 0.630 0.570 0.660
Heartbeat 0.751 0.751 0.756 0.756
MotorImagery 0.580 0.600 0.640 0.660
NATOPS 0.927 0.922 0.972 0.961
SelfRegulationSCP2 0.538 0.544 0.550 0.600
Table 2: Test acccuracy of augmenting Zz-Gril to TapNet and TodyNet on UEA Multivariate Time Series Classification Datasets.
Given a sample of multivariate time-series data with $m$ time-series, we splice each time-series into a sequence of time-series, each of length $w$. This splicing is done by a moving window of width $w$, where consecutive windows have an overlap of $\lambda$. We calculate the Pearson correlation coefficient Pearson & Lee (1903) between time-series in each window. We construct a graph for each window, where each time-series (of length $w$) is a node, and edges with the Pearson correlation values as weights connect them. Thus, we get a complete graph with $m$ nodes. We select the top $k$ percentile of these edges to build the final graph in each window. This way we obtain a sequence of graphs. The topological information of this sequence of graphs encodes the evolution of correlation between time series. Alternately, we can also track the evolving time-series by converting it into a sequence of point clouds. From the sequence of time-series described above, we consider each time-series of length $w$ as a point in $\mathbb{R}^{w}$. Thus, if we have $m$ time series, we have $m$ points in $\mathbb{R}^{w}$ in each window. Thus, we get a sequence of point clouds in $\mathbb{R}^{w}$. This sequence of point clouds tracks the evolution of each time-series, and a Vietoris-Rips like construction on this point cloud tracks the evolution of the interaction between time-series. Refer to Figure 6 for an illustration. Refer to Appendix LABEL:app:exp for hyperparameters.
5.2 UEA Multivariate Time Series Classification
UEA Multivariate Time Series Classification (MTSC) Bagnall et al. (2018) archive comprises of real-world multivariate time series data and is a widely recognized benchmark in time series analysis. The UEA MTSC collection encompasses a diverse range of application domains such as healthcare (ECG or EEG data), motion recognition (recorded using wearable sensors). See Table LABEL:tab:app:dataset_info in Appendix LABEL:app:exp.
The datasets are preprocessed and split into training and testing sets. For this set of experiments, we convert the multivariate time series into a sequence of point clouds and compute Zz-Gril. We choose 5 datasets which have at least 7 multivariate time series. This ensures that each point cloud, in the sequence of point clouds, has at least 7 points, giving meaningful topological information. Refer to Figure 7 for a visualization of Zz-Gril on FingerMovements dataset processed as a sequence of point clouds.
Figure 7: Heatmap of Zz-Gril on the FingerMovements dataset. Zz-Gril is computed at 36 center points, which is plotted as a 6x6 grid. The top row represents three samples with label 1 and the bottom row represents three samples with label 0. We can see the similarity in the Zz-Gril signatures between samples of the same class and clear differences between samples of different classees.
In Table 2, we compare the performance of augmenting Zz-Gril to two specifically tailored machine learning models for multivariate time series classification on UEA MTSC datasets, TapNet Zhang et al. (2020) and TodyNet Liu et al. (2024). We can see from the table that Zz-Gril is adding meaningful topological information on most datasets which is seen as an improvement in performance. We picked these two models for augmenting Zz-Gril because of two reasons: (i) availability and ease-of-use of codebase, (ii) these models already have good performance on UEA MTSC Datasets and we wanted to test the additional value of the topological information that Zz-Gril adds. In order to do a comprehensive evaluation of Zz-Gril framework, we compare Zz-Gril in isolation with other methods Li et al. (2021); Tang et al. (2022); Karim et al. (2019); Schäfer & Leser (2017) on UEA MTSC task. We report the results in Table 1. We see that Zz-Gril has comparable performance. Here, we would like to remind the reader that Zz-Gril is a general framework which does not need to be tailored for these datasets in particular. The aim for these experiments is to show that topological information is meaningful and is captured by Zz-Gril. Further, to show that we need both the spatial and the temporal information to be captured together, we compare Zz-Gril with standard zigzag persistence at various spatial scales. We report the results in Table 3. We can see that Zz-Gril performs better than standard zigzag persistence even across different scales.
Dataset Zigzag (scale = 0.5) Zigzag (scale = 0.7) Zizag (full graph) Zz-Gril
FingerMovements 0.490 0.490 0.530 0.590
Heartbeat - 0.682 0.687 0.721
MotorImagery 0.510 0.520 0.540 0.580
NATOPS - 0.805 0.816 0.850
SelfRegulationSCP2 0.483 0.489 0.500 0.522
Table 3: Accuracy of 1-parameter zigzag persistence at different scales compared with Zz-Gril. We use a Random Forest Classifier for classifying the topological signatures. The empty entries denote that the model does not train on the topological features because the features are not distinct enough at that scale.
We perform another set of experiments where we compare the performance of converting the multivariate time series into sequence of point clouds versus converting it into sequence of graphs for extracting topological information. We report the results in Table 5. We can see that both of these approaches are viable choices.
We report the computation times in Table 4 in. We can see from the table that Zz-Gril is, indeed, practical to use.
Dataset Time (Zz-Gril point clouds) Time (Zz-Gril graphs)
FingerMovements 569s 138s
NATOPS 413s 68s
SelfRegulationSCP2 153s 27s
Table 4: Computation times for Zz-Gril. The values represent the computation times for both training and testing splits. All the experiments were performed on Intel(R) Xeon(R) Gold 6248R CPU and NVIDIA Quattro RTX 6000 GPU.
Dataset TodyNet + Zz-Gril point clouds TodyNet + Zz-Gril graphs
FingerMovements 0.660 0.680
NATOPS 0.961 0.945
SelfRegulationSCP2 0.600 0.594
Table 5: Comparison of converting multivariate time series as a sequence of point clouds versus converting it as a sequence of graphs to extract the topological information.
5.3 Sleep Stage Classification
We use ISRUC-S3 dataset, which is a part of the ISRUC (Iberian Studies and Research on Sleep) Sleep Dataset Khalighi et al. (2016). ISRUC-S3 contains PSG recordings from 10 subjects. Each recording includes multiple physiological signals, such as: EEG, ECG. The dataset is annotated with sleep stage labels for each epoch (30 second window). There are 5 sleep stage labels: Wake (W), N1, N2, N3 (non-REM stages) and REM. We use STDP-GCN Zhao et al. (2023) as the machine learning model to augment. In this experiment, we convert the time series into sequence of graphs and compare the performance in Table 6. We can see that augmenting Zz-Gril increases both the accuracy and the overall F-1 score.
Methods Accuracy F-1 Overall F-1 Wake F-1 N-1 F-1 N2 F-1 N3 F-1 REM
SVM Alickovic & Subasi (2018) 73.3 72.1 86.8 52.3 69.9 78.6 73.1
RF Memar & Faradji (2017) 72.9 70.8 85.8 47.3 70.4 80.9 69.9
MLP+LSTM Dong et al. (2017) 77.9 75.8 86.0 46.9 76.0 87.5 82.8
CNN+BiLSTM Supratak et al. (2017) 78.8 77.9 88.7 60.2 74.6 85.8 80.2
CNN Chambon et al. (2018) 78.1 76.8 87.0 55.0 76.0 85.1 80.9
ARNN+RNN Phan et al. (2019) 78.9 76.3 83.6 43.9 79.3 87.9 86.7
STGCN Jia et al. (2020) 79.9 78.7 87.8 57.4 77.6 86.4 84.1
MSTGCN Jia et al. (2021) 82.1 80.8 89.4 59.6 80.6 89.0 85.6
STDP-GCN Zhao et al. (2023) 82.6 81.0 83.5 62.9 83.1 86.0 90.6
STDP-GCN + Zz-Gril 83.8 81.1 88.6 58.1 85.4 82.7 90.9
Table 6: Accuracy and F-1 scores of augmenting Zz-Gril to STDP-GCN and testing on ISRUC-S3 sleep classification dataset. We report the accuracy and F-1 scores for each model.
We would like to clarify that the aim, for both sets of experiments, is primarily to show that an increase in accuracy upon augmentation signifies that Zz-Gril captures meaningful topological information which can be used to improve the existing models.
6 Conclusion
In this paper, we proposed QZPH as a framework to capture both static and dynamic topological features in time-varying data. We proposed Zz-Gril, a stable and computationally efficient topological invariant to address the challenges of integrating MPH and ZPH. Through applications in various domains, including sleep-stage detection, we showed that augmenting machine learning models with Zz-Gril improves the performance. These results highlight the potential of integrating topological information to address complex challenges while analyzing time-evolving data.
Limitations. One of the limitations of this method, like most other persistent homology based methods, is its scalability. With the increasing size of the bifiltrations, the number of sampled center points to compute Zz-Gril needs to increase. Consequently, the number of zigzag persistence computations and the length of the zigzag filtrations over which they need to be performed increase. However, recent advances in persistent homology computations in the context of ML Nigmetov & Morozov (2024) give us the hope.
7 Acknowledgement
This work is partially supported by NSF grants DMS-2301360 and CCF-2437030. We acknowledge the discussion with Michael Lesnick who pointed out the theory about initial/terminal functors in the context of limits and colimits.
Appendix A Missing Definitions and Proofs
There is a notion of proximity on the space of zigzag modules in terms of the interleaving distance. In Botnan & Lesnick (2018), the authors define interleaving distance between two zigzag modules by including them into $\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$-indexed modules.
The interleaving distance on $\mathbb{R}^{n}$-indexed persistence modules is known and well-defined. We briefly recall the definition here. We refer the readers to Botnan & Lesnick (2018) for additional details.
Definition A.1 ($u$-shift functor).
The $u$-shift functor $(-)_{u}:\mathbf{vec}^{\mathbb{R}^{n}}\to\mathbf{vec}^{\mathbb{R}^{n}}$, for $u\in\mathbb{R}^{n}$, is defined as follows:
  1. 1.
    For $M\in\mathbf{vec}^{\mathbb{R}^{n}}$, $M_{u}$ is defined as $M_{u}(x)=M(x+u)$ for all $x\in\mathbb{R}^{n}$ and $M_{u}(x_{1}\leq x_{2})=M(x_{1}+u\leq x_{2}+u)$ for all $x_{1}\leq x_{2}\in\mathbb{R}^{n}$,
  2. 2.
    Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $F:M\to N$ be a morphism. Then, the corresponding morphism $F_{u}:M_{u}\to N_{u}$ is defnied as $F_{u}(x)=F(x+u):M_{u}(x)\to N_{u}(x)$ for all $x\in\mathbb{R}^{n}$.
Definition A.2 ($\epsilon$-interleaving).
Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $\epsilon\in[0,\infty)$ be given. We will denote $(-)_{\bm{\epsilon}}$ to be the shift functor corresponding to the vector $\bm{\epsilon}=\epsilon(1,1,\ldots,1)$. We say $M$ and $N$ are $\epsilon$-interleaved if there are natural transformations $F:M\to N_{\bm{\epsilon}}$ and $G:N\to M_{\bm{\epsilon}}$ such that
  1. 1.
    $G_{\bm{\epsilon}}\circ F=\varphi_{M}^{2\bm{\epsilon}}$,
  2. 2.
    $F_{\bm{\epsilon}}\circ G=\varphi_{N}^{2\bm{\epsilon}}$,
where $\varphi_{M}^{u}:M\to M_{u}$ is the natural transformation whose restriction to each $M(x)$ is the linear map $M(x\leq x+u)$ for all $x\in\mathbb{R}^{n}$.
Definition A.3 (Interleaving distance).
Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. The interleaving distance $d_{\mathcal{I}}(M,N)$ between $M$ and $N$ is defined as
[…]
and $d_{\mathcal{I}}(M,N)=\infty$ if there exists no interleaving.
We need the following two definitions to define interleaving distance between two zigzag modules.
Definition A.4 (Left Kan Extension).
Let $P$ and $Q$ be two posets. Let $F:P\to Q$ be a functor. Let $P[F\leq q]$ denote the set $P[F\leq q]\coloneqq\{p\in P:F(p)\leq q\}$. Given a persistence module $M:P\to\mathbf{vec}$, the left Kan extension of $M$ along $F$ is a functor $\text{Lan}_{F}(M):Q\to\mathbf{vec}$ given by
$\text{Lan}_{F}(M)(q)\coloneqq\mathsf{colim}M|_{P[F\leq q]},$
along with internal morphisms given by the universality of colimits.
Definition A.5 (Block Extension Functor Botnan & Lesnick (2018)).
Let $\mathbb{U}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the poset $\mathbb{U}\coloneqq\{(a,b):a\leq b\}$. Let $\mathfrak{i}:\mathbb{ZZ}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the inclusion. Let $(-)|_{\mathbb{U}}:\mathbf{vec}^{\mathbb{R}^{\mathsf{op}}\times\mathbb{R}}\to\mathbf{vec}^{\mathbb{U}}$ denote the restriction. Then, the block extension functor $E:\mathbf{vec}^{\mathbb{ZZ}}\to\mathbf{vec}^{\mathbb{U}}$ is defined as
$E\coloneqq(-)|_{\mathbb{U}}\circ\text{Lan}_{\mathfrak{i}}(\circ).$
Definition A.6 (Interleaving distance on zigzag modules).
Let $M,N$ be two zigzag modules. Then
$d_{\mathcal{I}}(M,N)\coloneqq d_{\mathcal{I}}(E(M),E(N)).$
We extend this definition of interleaving distance to quasi zigzag persistence modules. Let $\mathbb{UU}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ be the analog of $\mathbb{U}$, i.e, $\mathbb{UU}\coloneqq\{(a,b,c):a\leq b\}$. Let $\tilde{\mathfrak{i}}:\mathbb{ZZ}\times\mathbb{Z}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ denote the inclusion. Then, define $\tilde{E}\coloneqq(-)|_{\mathbb{UU}}\circ\text{Lan}_{\tilde{\mathfrak{i}}}(\circ)$, analogous to the block extension functor.
Definition A.7.
Let $M,N$ be two quasi zigzag persistence modules. Then, the interleaving distance between $M$ and $N$ is given by
$d_{\mathcal{I}}(M,N)\coloneqq d_{\mathcal{I}}(\tilde{E}(M),\tilde{E}(N)).$
Lemma A.8.
$\tilde{E}$ sends interval modules on $\mathbb{ZZ}\times\mathbb{Z}$ to block interval modules.
This is analogous to Lemma 4.1 in Botnan & Lesnick (2018).
Lemma A.9.
Let $M$ be a p.f.d quasi zigzag persistence module. If $M\cong\bigoplus_{k\in K}I_{k}$, then $\tilde{E}(M)\cong\bigoplus_{k\in K}\tilde{E}(I_{k})$, where $K$ is an indexing set.
Proof.
The functor $\text{Lan}_{\tilde{\mathfrak{i}}}$ preserves direct sums MacLane (1971), as it is left-adjoint to the canonical restriction functor. The restriction functor also preserves direct sums. These two facts combined with Lemma A.8 give the result. ∎
Let $I$ denote a subposet in $\mathbb{ZZ}\times\mathbb{Z}$. For ϵ+italic-$\epsilon\in\mathbb{Z}_{+}$, the $\epsilon$-thickening of $I$ is defined as
$I^{\epsilon}\coloneqq\left\{\mathbf{r}\in\mathbb{ZZ}\times\mathbb{Z}:\exists\mathbf{q}\in I\text{ such that }||\mathbf{r}-\mathbf{q}||_{\infty}\leq\epsilon\right\}.$
Recall that $(\mathbb{ZZ}\times\mathbb{Z})$ denotes the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals.
Definition A.10.
Let $M$ and $N$ be two quasi zigzag persistence modules. The erosion distance is defined as:
$d_{\mathcal{E}}(M,N)\coloneqq\inf\{\epsilon\geq 0:\forall I\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\mathsf{rk}^{M}(I)\geq\mathsf{rk}^{N}(I^{\epsilon})\text{ and }\mathsf{rk}^{N}(I)\geq\mathsf{rk}^{M}(I^{\epsilon})\}.$
Proposition 3.4.
Given two quasi zigzag persistence modules $M$ and $N$, $d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{I}}(M,N)$ where $d_{\mathcal{I}}$ denotes the interleaving distance between $M$ and $N$.